# Orbital paramagnetism without density of states enhancement in nodal-line semimetal ZrSiS

### v1

@@ Abstract
Unconventional orbital paramagnetism without enhancement of the density of states was recently discovered in the nodal-line semimetal ZrSiS. Here, we propose a novel interband mechanism of orbital paramagnetism associated with the negative curvature of energy dispersions, which successfully explains the observed anomalous orbital paramagnetism. This negative curvature arises from energy fluctuations along the nodal line, inherent in realistic nodal-line materials. Our new mechanism indicates that such orbital paramagnetism serves as strong evidence for the existence of nodal lines not only in ZrSiS but potentially in various other nodal-line materials as well.
@@
Orbital magnetism is one of the fundamental properties of solids, rooted in the seminal research of Landau and Peierls on free electron and tight-binding models [1, 2]. In Dirac electron systems such as bismuth and graphene, significant orbital diamagnetism arises from the interband effect of a magnetic field [3, 4, 5, 6, 7]. This discovery has shown that orbital magnetism is highly sensitive to the band structure of crystals, leading to extensive research into their relationship that has continued to this day. Orbital magnetism is also a bulk property that probes the unique energy dispersions and wave functions of electronic states with nontrivial topology, such as those in Weyl semimetals [3, 7, 8, 4, 5, 9, 9, 10, 11, 6, 12, 13] and topological insulators [14, 15, 16, 17]. We expect orbital magnetism to also characterize nodal-line semimetals [18, 19, 20, 21, 22, 23, 24], which are usually confirmed through a combination of angle-resolved photoemission spectroscopy [25, 26, 27], quantum oscillations [28, 29, 30], and transport phenomena [23, 24, 7, 31, 32, 33, 34, 35].
Recently, unconventional paramagnetism was observed in the nodal-line semimetal ZrSiS at low temperatures when a magnetic field is applied along the $C_{4}$ rotation axis [30]. This paramagnetism is not explained by the Pauli paramagnetism estimated from the observed low density of states (DOS). Additionally, when the magnetic field is applied perpendicular to the $C_{4}$ axis, ZrSiS exhibits diamagnetism, showing strong anisotropy. These results strongly suggest that the unconventional behaviors of ZrSiS originate from the orbital effect rather than the spin effect. A similar paramagnetism has been observed in another nodal-line semimetal SrAs3 [36]. Furthermore, as the temperature increases, the paramagnetism along the $C_{4}$ axis of ZrSiS decreases and eventually changes to diamagnetism at around 120K, which is also an unexpected behavior. Despite several efforts to understand these anomalous behaviors, their microscopic origins, probably due to the nodal line [30, 37], remain elusive.
@@
At present, a few origins of orbital paramagnetism are known, including the Van Hove singularity [38, 39], flatband [40, 41], and other mechanisms [37, 42]. In most of these cases, however, orbital paramagnetism is accompanied by a strong enhancement of a DOS, and thus it is concealed by the large Pauli paramagnetism. In this Letter, we propose a new mechanism to explain the orbital paramagnetism without an enhanced DOS. We first present a quantitative analysis of the orbital magnetic susceptibility in ZrSiS using the density functional theory (DFT) calculations and the effective models based on them, successfully illustrating the observed paramagnetism, temperature dependence, and anisotropy. To further understand the origin of this orbital paramagnetism, we derive a simple effective model. The analysis of it shows that the observed orbital paramagnetism without DOS enhancement is attributed to the interband effect between two energy dispersions with negative curvature. This negative curvature originates from the energy fluctuation along the nodal line, which is inherent in realistic nodal-line materials. This novel mechanism indicates that such orbital paramagnetism serves as strong evidence for the nodal lines not only in ZrSiS but potentially in various other nodal-line materials as well.
@@
Figure 1: Positions of the nodal points in the Brillouin zone. The dimensionless wave numbers $\bar{k}_{x}=k_{x}a/(2\pi)$, $\bar{k}_{y}=k_{y}a/(2\pi)$, and $\bar{k}_{z}=k_{z}c/(2\pi)$ are shown. (a) Color indicates the energy of the middle of the gap $(E_{c}({\bm{k}})+E_{v}({\bm{k}}))/2-E_{F}$ with $E_{c}({\bm{k}})$ and $E_{v}({\bm{k}})$ being the energy of the bottom (top) of the conduction (valence) band at $\bm{k}$, and $E_{\rm F}$ is the Fermi energy. (b) Color indicates the magnitude of the gap due to the SOI, $E_{c}({\bm{k}})-E_{v}({\bm{k}})$. (c) Energy dispersions near $E_{\rm F}$ along the nodal lines (i) and (ii). The valence (conduction) band is shown in solid (dashed) line.
ZrSiS has a $P4/mnn$ structure [43, 44] and has a set of nodal lines [45, 30, 46, 47]. To capture the characteristics of these nodal lines in detail, we perform DFT calculations using Quantum-ESPRESSO and Wannier90 packages [48, 49, 50, 51, 52, 53] with the lattice parameters of $a=3.55$ and $c=8.07$ [54]. The details of this calculation are shown in Supplemental Material (SM) [55]. Figure 1 shows the positions of the nodal points in the Brillouin zone, at which the bottom of the conduction band ($E_{c}({\bm{k}})$) and the top of the valence band ($E_{v}(\bm{k})$) are close to each other. The colors in Fig. 1(a) indicate the energy of the middle of the gap relative to the Fermi energy $E_{F}$, i.e., $(E_{c}(\bm{k})+E_{v}(\bm{k}))/2-E_{F}$, and those in Fig. 1 (b) indicate the magnitude of the gap $E_{c}(\bm{k})-E_{v}(\bm{k})$ due to the spin-orbit interaction (SOI). Four nodal lines (i) – (iv) are indicated, which are in the planes of $k_{z}=\pi/c$, $k_{z}=0$, $|k_{x}|=|k_{y}|$, and $k_{x}=0$ or $k_{y}=0$, respectively. Note that the nodal line protected by the nonsymmorphic symmetry [45] is omitted because it is approximately 1eV away from $E_{F}$. From Fig. 1(a), we can see that the nodal lines (i) and (ii) form closed lines in the Brillouin zone on the $k_{z}c=\pi$ and $k_{z}c=0$ planes, respectively, both of which cross $E_{F}$. As shown in Fig. 1 (c), the energy deviations around $E_{F}$ are about ±$\pm 0.25$ eV for (i) and ±$\pm 0.05$ eV for (ii), which play important roles in the later calculations. The nodal lines (iii) and (iv) form one-dimensional lines along the $k_{z}$ direction. While the nodal line (iii) is located close to $E_{F}$, (iv) is away from $E_{F}$, so that we neglect the nodal line (iv) in the following.
@@
where X=i or ii is the index for the nodal lines and $\sigma_{0}$, $\sigma_{x}$, and $\sigma_{z}$ are the identity and Pauli matrices. The first term represents the two parabolic bands with masses of different signs. As shown in Fig. 2, the overlap of the bands is $2\Delta$. The second term in Eq. (1) hybridizes the two bands where $v_{z}$ represents the velocity in the $z$ direction. The ${\bm{k}}$ points on the nodal line satisfies $\frac{\hbar^{2}}{2m^{*}}(k_{x}^{2}+k_{y}^{2})-\Delta=0$ and $k_{z}=0$, indicating that the nodal line forms a circle in the $k_{x}$-$k_{y}$ planes at $E_{F}=0$. The last term gives the fluctuation of energy along the nodal line, and the energy deviation is ±$\pm 2\eta^{\text{(X)}}k_{\rm R}^{2}$ with $k_{\rm R}=\sqrt{2m^{*}\Delta}/\hbar$ being the radius of the nodal line. The similar energy fluctuation has been discussed in Refs. [37, 56, 57, 58, 31]. In the following, we omit the superscript of $\eta$ when distinction is not necessary.
Figure 2: Energy dispersions and DOS for model (1) at $k_{z}=0$ with $\tilde{\eta}=\eta k_{\rm R}^{2}/\Delta=0.2$. The red line indicates the nodal line. $\tilde{k}_{x}=k_{x}/k_{\rm R}$ and $\tilde{k}_{y}=k_{y}/k_{\rm R}$ are dimensionless wave numbers. The symbol $\bm{\times}$ indicates the nodal point at which the tangent of the nodal line is parallel to $\tilde{k}_{x}$.
Note that the effective model of Eq. (1) does not have the gap. As discussed later, the effect of gap is negligible when we compare the results with the experiment.
@@
where ${\mathcal{G}}=\mathcal{G}({\bm{k}},i\varepsilon_{n})=[i\varepsilon_{n}-\mathcal{H}^{\text{(X)}}+\mu]^{-1}$ is the thermal Green’s function, $\varepsilon_{n}=(2n+1)\pi k_{\rm B}T$ ($n\in\mathbb{Z})$ is the Matsubara frequency, and $\gamma_{i}$ is the velocity operator of in the $i$ ($i=x,y,z$) direction defined by $\gamma_{i}=\frac{\partial{\mathcal{H}}^{\text{(X)}}}{\partial k_{i}}$. This formula has included the spin degeneracy. The orbital magnetic susceptibility in other directions is obtained by a cyclic replacement of $x$, $y$, and $z$. We numerically evaluate Eq. (3) for the model (1), employing the quasi-Monte Carlo method [60, 61, 62, 63, 64] for ${\bm{k}}$ summation, and the sparse-ir method [65, 66, 67] for Matsubara summation. The cutoff for ${\bm{k}}$ integral is set to $\Lambda_{\rm R}=1000k_{\rm R}$ in the radial direction and $\Lambda_{z}=(l_{z})^{-1}$ in the $z$ direction with $l_{z}=\hbar v_{z}/\Delta$. The results for $\chi_{z}$ and $\chi_{x}$ are shown in Fig. 3 (a) and (b), respectively, for several values of $\tilde{\eta}=\eta k_{\rm R}^{2}/\Delta$. For the case without the energy fluctuation ($\tilde{\eta}=0$), $\chi_{z}$ is diamagnetic for every $\mu$ while $\chi_{x}$ has a sharp peak at $\mu=0$ originating from the interband effect of the 2D Dirac electrons [6]. It is confirmed that our calculation reproduces the previous results [24] in the limit of $\Lambda_{\rm R}\to\infty$ and $\Lambda_{z}\to\infty$ [see SM[55] for detail].
@@
Finite $\eta$ gives a drastic effect on $\chi_{z}$: $\chi_{z}$ has a broad peak around $\mu=0$, whose width is approximately $2\eta k_{\rm R}^{2}$. Furthermore, for $\tilde{\eta}>0.1$, the value at the peak is positive, meaning orbital paramagnetism. The inset of Fig. 3 (a) shows the Landau-Peierls (LP) contribution [1, 2, 39, 68], or the intraband contribution, which is negative for all the value of $\mu$. Therefore, we conclude that the obtained orbital paramagnetism near $\mu=0$ is due to an interband effect. As we noted before, the orbital paramagnetism is usually accompanied by a large DOS, while the present result suggests an interband orbital paramagnetism without the enhancement of the DOS. Note that the DOS near $\mu=0$ is small as shown in Fig. 2. The mechanism is discussed in detail later.
@@
Figure 3: (a) Orbital magnetic susceptibility $\chi_{z}$ as a function of chemical potential when $k_{B}T=0.01\Delta$ for several values of $\tilde{\eta}=\eta k_{\rm R}^{2}/\Delta$. (b) Orbital magnetic susceptibility $\chi_{x}$ when $k_{B}T=0.05\Delta$. They are normalized by $\chi_{0}\equiv\frac{e^{2}}{\hbar^{2}}Vl_{z}\Delta$ with $V$ being the volume and $l_{z}=\hbar v_{z}/\Delta$. Inset: LP contribution for $\tilde{\eta}=0.3$ and $k_{B}T=0.1\Delta$.
@@
As shown in Fig. 3 (b), $\chi_{x}$ is not an even function of the chemical potential for finite $\eta$. We find the relations $\chi_{x}(\mu,\eta)=\chi_{x}(-\mu,-\eta)$ and $\chi_{x}(\mu,\eta)=\chi_{y}(-\mu,\eta)$ from the analytical expressions, shown in SM [55]. This behavior is understood as follows. At each point on the nodal line, a Dirac dispersion ±$\propto\pm\sqrt{q_{x}^{2}+q_{y}^{2}}$ is formed, where $q_{x}$-$q_{y}$ plane is perpendicular to the tangent of the nodal line at the point. Therefore, when the magnetic field is parallel to the tangent of the nodal line, the delta function–like orbital diamagnetism ($\propto-\delta(\mu-\varepsilon_{0}))$ [59, 6] is negatively maximized with $\varepsilon_{0}$ being the energy of the Dirac point. As we can see from the symbol $\times$ in Fig. 2, when $\eta$ is positive, the energy of the nodal point at which the tangent of the nodal line is parallel to $\tilde{k}_{x}$ is negative, i.e., $\varepsilon_{0}<0$. Thus, $\chi_{x}$ takes the negatively maximal value at $\mu=\varepsilon_{0}<0$. The magnitude of $\chi_{x}$ becomes smaller as $\eta$ increases, since the region of the nodal line parallel to $\tilde{k}_{x}$ becomes smaller.
@@
Comparison with experiments.— We evaluate the orbital magnetic susceptibility for ZrSiS using the parameters obtained by the DFT calculations. For $\chi_{z}$, we consider $\chi_{z}^{\rm tot}=\chi_{z}^{\text{(i)}}+\chi_{z}^{\text{(ii)}}+\chi_{z}^{\text{(iii)}}+\chi_{\rm Pauli}$, where $\chi_{z}^{\text{(X)}}$ represents the contribution from the nodal line X (X=i, ii, iii), and the Pauli paramagnetism $\chi_{\rm Pauli}=0.125\times 10^{-4}$ emu/mol [30] is included. We evaluate $\chi_{z}^{\text{(i)}}$ and $\chi_{z}^{\text{(ii)}}$ from Eq. (1). $\chi_{z}^{\text{(iii)}}$ is obtained from the model (2) by the $k_{z}$ average of the known result for the 2D Dirac electron system [39]
@@
[…]
@@
From the DFT calculations, we obtain $\tilde{\eta}^{\text{(i)}}=0.31$ for (i), $\tilde{\eta}^{\text{(ii)}}=0.06$ for (ii), and $\Delta_{\text{SOI}}=15$ meV and $\varepsilon_{0}(k_{z})=(-0.025-0.07|k_{z}|c/\pi)$ eV for (iii), respectively. The other parameters are adjusted to fit the experimental data [30], which are $\Delta=0.8$ eV, $m^{*}/m_{0}=0.12$, $v_{z}/c_{0}=4\times 10^{-4}$, and $v_{\rm Dirac}/c_{0}=9\times 10^{-4}$ with $m_{0}$ and $c_{0}$ being the bare electron mass and the speed of light in vacuum, respectively. Figure 4 shows the result of $\chi_{z}^{\text{tot}}$ and $\chi_{x}^{\text{tot}}$, which are in good agreement with the experiment.

### v2

@@ Abstract
Unconventional orbital paramagnetism without enhanced density of states was recently discovered in the nodal-line semimetal ZrSiS. We propose a novel interband mechanism, linked to the negative curvature of energy dispersions, which successfully accounts for the observed anomalous response. This negative curvature originates from energy variation along the nodal line, inherent in realistic nodal-line materials. Our results suggest that such orbital paramagnetism provides strong evidence for the presence of nodal lines in ZrSiS, and serves as a hallmark of other nodal-line materials.
@@
Orbital magnetism is one of the fundamental properties of solids, rooted in the seminal work of Landau and Peierls on free-electron and tight-binding models [1, 2]. In Dirac electron systems such as bismuth and graphene, significant orbital diamagnetism arises from the interband effect of a magnetic field [3, 4, 5, 6, 7]. This discovery has shown that orbital magnetism is highly sensitive to the electronic structure, leading to extensive ongoing research into their relationship. Orbital magnetism is a bulk response that is also sensitive to the geometric properties encoded in the wave functions, such as nontrivial topology and quantum geometry. Weyl semimetals [3, 7, 8, 4, 5, 9, 9, 10, 11, 6, 12, 13], topological insulators [14, 15, 16, 17], and nodal-line semimetals [18, 19, 20, 21, 22, 23, 24, 7, 25] are prominent examples of these connections. Among these, nodal-line semimetals have been less characterized experimentally. Consequently, orbital magnetism is expected to offer a more direct, bulk-sensitive probe reflecting their characteristic electronic structures, complementing standard techniques such as angle-resolved photoemission spectroscopy [26, 27, 28], quantum oscillations [29, 30, 31], and transport studies [32, 33, 34, 35].
Orbital paramagnetism, as an unconventional orbital response, has recently been observed in the nodal-line semimetal ZrSiS at low temperatures, when a magnetic field is applied along the $C_{4}$ rotation axis [31]. This finding challenges the established understanding based on Larmor’s classical picture and Landau’s theory, both of which regard orbital magnetism as inherently diamagnetic. Furthermore, as the temperature increases, the paramagnetism decreases and eventually changes to diamagnetism at around 120 K, which is also an unexpected behavior. A similar paramagnetism has also been observed in another nodal-line semimetal, SrAs3 [36]. These observations in nodal-line semimetals suggest a possible connection between the presence of nodal lines and the emergence of such orbital paramagnetism, though its microscopic origin remains elusive.
@@
At present, a few origins of orbital paramagnetism are known, including the Van Hove singularity [37, 38, 39], flatband [40, 41], and other mechanisms [42, 43]. In most of these cases, however, orbital paramagnetism is accompanied by a strong enhancement of density of states (DOS), and thus it is often masked by the large Pauli paramagnetism. In this Letter, we propose a new mechanism to explain the orbital paramagnetism without enhanced DOS. We first present a quantitative analysis of the orbital magnetic susceptibility in ZrSiS using the effective models based on the density functional theory (DFT) calculations, successfully elucidating the observed paramagnetism, temperature dependence, and anisotropic behaviors [31, 44]. To further understand the origin of this orbital paramagnetism, we derive a simple effective model. Our results show that the observed orbital paramagnetism without DOS enhancement is attributed to the interband effect associated with negative curvature in the energy dispersion. This negative curvature originates from the energy variation along the nodal line, which is inherent in realistic nodal-line materials. While dispersions with negative curvature can in principle occur in other systems, they emerge more naturally and systematically in nodal-line semimetals due to the presence of line-like band crossing. This novel mechanism suggests that such orbital paramagnetism provides strong evidence for the nodal lines not only in ZrSiS but also potentially in various other nodal-line materials.
@@
Figure 1: Positions of the nodal points in the Brillouin zone. The dimensionless wave numbers $\bar{k}_{x}=k_{x}a/(2\pi)$, $\bar{k}_{y}=k_{y}a/(2\pi)$, and $\bar{k}_{z}=k_{z}c/(2\pi)$ are shown. (a) Color indicates the energy of the middle of the gap $(E_{c}({\bm{k}})+E_{v}({\bm{k}}))/2-E_{F}$ with $E_{c}({\bm{k}})$ and $E_{v}({\bm{k}})$ being the energy of the bottom (top) of the conduction (valence) band at $\bm{k}$, and $E_{\rm F}$ is the Fermi energy. (b) Color indicates the magnitude of the gap due to the SOI, $E_{c}({\bm{k}})-E_{v}({\bm{k}})$. (c, d) Energy dispersions near $E_{\rm F}$ (c) along the nodal lines (i) and (ii), and (d) along the nodal line (iii). The valence (conduction) band is shown in solid (dashed) line.
ZrSiS has a $P4/nmm$ structure [45, 46] and has a set of nodal lines [47, 31, 48, 49]. To capture the characteristics of these nodal lines in detail, we perform DFT calculations using Quantum-ESPRESSO and Wannier90 packages [50, 51, 52, 53, 54, 55] with the lattice parameters of $a=3.55$ Å and $c=8.07$ Å [56]. The details of these calculations are shown in Supplemental Material (SM) [57]. Figure 1 shows the positions of the nodal points in the Brillouin zone, at which the bottom of the conduction band ($E_{c}({\bm{k}})$) and the top of the valence band ($E_{v}(\bm{k})$) are close to each other. The colors in Fig. 1(a) indicate the energy of the middle of the gap relative to the Fermi energy $E_{F}$, i.e., $(E_{c}(\bm{k})+E_{v}(\bm{k}))/2-E_{F}$, and those in Fig. 1 (b) indicate the magnitude of the gap $E_{c}(\bm{k})-E_{v}(\bm{k})$ due to the spin-orbit interaction (SOI). Four nodal lines (i) – (iv) are indicated, which are in the planes of $k_{z}=\pi/c$, $k_{z}=0$, $|k_{x}|=|k_{y}|$, and $k_{x}=0$ or $k_{y}=0$, respectively. Note that the nodal line protected by the nonsymmorphic symmetry [47] is omitted because it is approximately 1eV away from $E_{F}$. From Fig. 1(a), we can see that the nodal lines (i) and (ii) form closed lines in the Brillouin zone on the $k_{z}c=\pi$ and $k_{z}c=0$ planes, respectively, both of which cross $E_{F}$. As shown in Fig. 1 (c), the energy deviations around $E_{F}$ are approximately $\pm 0.25$ eV for (i) and $\pm 0.05$ eV for (ii), which play important roles in the later calculations. Figure 1(d) represents the energy of the valence and conduction bands along the nodal line (iii), where the Fermi level lies outside the Dirac mass gap. Nodal lines (iii) and (iv) form one-dimensional lines along the $k_{z}$ direction. While the nodal line (iii) is located close to $E_{F}$, (iv) is away from $E_{F}$, so that we neglect the nodal line (iv) in the following.
@@
where X=i or ii is the index for the nodal lines and $\sigma_{0}$, $\sigma_{x}$, and $\sigma_{z}$ are the identity and Pauli matrices. We have assumed a circular nodal line instead of the expected square nodal line for simplicity. This will be justified later by a comparison with the experimental data. The first term represents the two parabolic bands with masses of different signs. As shown in Fig. 2, the overlap of the bands is $2\Delta$. The second term in Eq. (1) hybridizes the two bands where $v_{z}$ represents the velocity in the $z$ direction. The $\bm{k}$} points on the nodal line satisfy $\frac{\hbar^{2}}{2m^{*}}(k_{x}^{2}+k_{y}^{2})-\Delta=0$ and $k_{z}=0$, indicating that the nodal line forms a circle in the $k_{x}$-$k_{y}$ planes at $E_{F}=0$. Finite $k_{z}$ opens gaps along the nodal line. The last term introduces variation of energy along the nodal line, and the energy deviation is $\pm 2\eta^{\text{(X)}}k_{\rm R}^{2}$ with $k_{\rm R}=\sqrt{2m^{*}\Delta}/\hbar$ being the radius of the nodal line. We will see below that this variation play a crucial role for the emergence of orbital paramagnetism. Similar energy variation has been discussed in Refs. [42, 58, 59, 60, 25]. In the following, we omit the superscript of $\eta$ when a distinction is not necessary. The DOS is also shown in Fig. 2 for $\tilde{\eta}=0$ and $0.2$, where $\tilde{\eta}$ is a dimensionless parameter $\tilde{\eta}=\eta k_{\rm R}^{2}/\Delta$. The energy variation results in the emergence of DOS around $E=E_{F}$. This variation also gives negative Gauss curvature of the $E$-$k_{x}$-$k_{y}$ surface at $(k_{x},k_{y},k_{z})=(0,\pm k_{\rm R},k_{z0})$ for the valence band and at $(\pm k_{\rm R},0,k_{z0})$ for the conduction band for $k_{z0}\neq 0$, i.e., in the presence of gaps. The curvature is illustrated in detail in SM [57].
[…]
Note that the effective model of Eq. (1) does not include the SOI effect. As discussed later, SOI hardly affects our results.
@@
where ${\mathcal{G}}=\mathcal{G}({\bm{k}},i\varepsilon_{n})=[i\varepsilon_{n}-\mathcal{H}^{\text{(X)}}+\mu]^{-1}$ is the thermal Green’s function, $\varepsilon_{n}=(2n+1)\pi k_{\rm B}T$ ($n\in\mathbb{Z})$ is the Matsubara frequency, and $\gamma_{i}$ is the velocity operator of in the $i$ ($i=x,y,z$) direction defined by $\gamma_{i}=\frac{\partial{\mathcal{H}}^{\text{(X)}}}{\partial k_{i}}$. Equation (3) is a general formula for noninteracting systems and incorporates the spin degeneracy. The orbital magnetic susceptibilities in other directions, $\chi_{x}$ and $\chi_{y}$, are obtained by a cyclic replacement of $x$, $y$, and $z$. We numerically evaluate Eq. (3) for the model Eq. (1), employing the quasi-Monte Carlo method [62, 63, 64, 65, 66] for $\bm{k}$} summation, and the sparse-ir method [67, 68, 69] for Matsubara summation. The momentum cutoffs are $\Lambda_{R}=1000k_{R}$ in the radial direction and $\Lambda_{z}=(l_{z})^{-1}$ in the $k_{z}$ direction with $l_{z}=\hbar v_{z}/\Delta$. The Matsubara summation includes a sufficient number of frequencies corresponding to the energy cutoff $\varepsilon_{\rm cut}=100\Delta$  [^1] In our approach, we adopt a finite but sufficiently large cutoff for Matsubara summation. Correspondingly, we choose the $\bm{k}$} integral cutoff to be much larger than the Matsubara frequency cutoff. This method confirms that the numerical integration can be carried out efficiently and with sufficient accuracy, when compared with the analytical results obtained in Ref. [24] and SM [57].
The results for $\chi_{z}$ and $\chi_{x}$ are shown in Fig. 3 (a) and (b), respectively, for several values of $\tilde{\eta}$. For the case without the energy variation ($\tilde{\eta}=0$), $\chi_{z}$ is diamagnetic for every $\mu$ while $\chi_{x}$ has a sharp peak at $\mu=0$ originating from the interband effect of the 2D Dirac electrons [6]. We confirmed that our calculation reproduces the previous results [24] in the limit of $\Lambda_{\rm R}\to\infty$ and $\Lambda_{z}\to\infty$ [see SM[57] for detail].
@@
Finite $\eta$ gives a significant effect on $\chi_{z}$: $\chi_{z}$ has a broad peak around $\mu=0$, whose width is approximately $2\eta k_{\rm R}^{2}$. Furthermore, for $\tilde{\eta}>0.1$, the value at the peak is positive, meaning orbital paramagnetism. The inset of Fig. 3 (a) shows the Landau-Peierls (LP) contribution [1, 2, 38, 71], or the intraband contribution, which is negative for all values of $\mu$. Therefore, we conclude that the obtained orbital paramagnetism near $\mu=0$ is due to an interband effect. As we noted before, while the orbital paramagnetism is usually accompanied by large DOS, the present result suggests an interband orbital paramagnetism without the enhancement of the DOS. The presence of two energetically close bands involving negative curvature is a crucial factor in inducing orbital paramagnetism in this system. While such a configuration can, in principle, arise in other materials, it remains primarily associated with nodal-line semimetals and is seldom realized elsewhere. The mechanism for this paramagnetism is discussed in detail later.
@@
Figure 3: (a) Orbital magnetic susceptibility $\chi_{z}$ as a function of chemical potential when $k_{B}T=0.01\Delta$ for several values of $\tilde{\eta}=\eta k_{\rm R}^{2}/\Delta$. (b) Orbital magnetic susceptibility $\chi_{x}$ when $k_{B}T=0.05\Delta$. They are normalized by $\chi_{0}\equiv\frac{e^{2}}{\hbar^{2}}Vl_{z}\Delta$ with $V$ being the volume and $l_{z}=\hbar v_{z}/\Delta$. Inset: LP contribution for $\tilde{\eta}=0.2$ and $k_{B}T=0.1\Delta$.
@@
As shown in Fig. 3 (b), $\chi_{x}$ is not an even function of the chemical potential for finite $\eta$. We find the relations $\chi_{x}(\mu,\eta)=\chi_{x}(-\mu,-\eta)$ and $\chi_{x}(\mu,\eta)=\chi_{y}(-\mu,\eta)$ from the analytical expressions, shown in SM [57]. This behavior is understood as follows. At each point on the nodal line, a Dirac dispersion $\propto\pm\sqrt{q_{x}^{2}+q_{y}^{2}}$ is formed, where $q_{x}$-$q_{y}$ plane is perpendicular to the tangent of the nodal line at the point. Therefore, when the magnetic field is parallel to the tangent of the nodal line, the delta function–like orbital diamagnetism ($\propto-\delta(\mu-\varepsilon_{0}))$ [61, 6] is negatively maximized with $\varepsilon_{0}$ being the energy of the Dirac point. As we can see from the symbol $\times$ in Fig. 2, when $\eta$ is positive, the energy of the nodal point at which the tangent of the nodal line is parallel to $\tilde{k}_{x}$ is negative, i.e., $\varepsilon_{0}<0$. Thus, $\chi_{x}$ takes the negatively maximal value at $\mu=\varepsilon_{0}<0$. The magnitude of $\chi_{x}$ becomes smaller as $\eta$ increases, since the region of the nodal line parallel to $\tilde{k}_{x}$ becomes smaller.
@@
Comparison with experiments.— We evaluate the orbital magnetic susceptibility for ZrSiS using the parameters obtained by the DFT calculations. For $\chi_{z}$, we consider $\chi_{z}^{\rm tot}=\chi_{z}^{\text{(i)}}+\chi_{z}^{\text{(ii)}}+\chi_{z}^{\text{(iii)}}+\chi_{\rm Pauli}$, where $\chi_{z}^{\text{(X)}}$ represents the contribution from the nodal line X (X=i, ii, iii), and the Pauli paramagnetism $\chi_{\rm Pauli}=0.125\times 10^{-4}$ emu/mol [31] is included. We evaluate $\chi_{z}^{\text{(i)}}$ and $\chi_{z}^{\text{(ii)}}$ from Eq. (1). $\chi_{z}^{\text{(iii)}}$ is obtained from the model Eq. (2) by the $k_{z}$ average of the known result for the 2D Dirac electron system [38]
@@
as $\chi_{z}^{\text{(iii)}}=\frac{V}{2\pi}\int dk_{z}\chi_{2D}(k_{z})$ where $f(\varepsilon,\mu)$ is the Fermi distribution function.
@@
From the DFT calculations [see Fig. 1], we obtain $\tilde{\eta}^{\text{(i)}}=0.31$ for (i), $\tilde{\eta}^{\text{(ii)}}=0.06$ for (ii), and $\Delta_{\text{SOI}}=15$ meV and $\varepsilon_{0}(k_{z})=(-0.025-0.07|k_{z}|c/\pi)$ eV for (iii), respectively. Note that these parameters ensure that the Fermi level lies outside the Dirac mass gap along (iii). The other parameters are adjusted to fit the experimental data [31], which are $\Delta=0.8$ eV, $m^{*}/m_{0}=0.12$, $v_{z}/c_{0}=4\times 10^{-4}$, and $v_{\rm Dirac}/c_{0}=9\times 10^{-4}$ with $m_{0}$ and $c_{0}$ being the bare electron mass and the speed of light in vacuum, respectively. Figure 4 shows the results of $\chi_{z}^{\text{tot}}$ and $\chi_{x}^{\text{tot}}$, where $\chi_{z}^{\rm tot}$ agrees well with the experimental data quantitatively and $\chi_{x}^{\rm tot}$ captures the observed trends qualitatively.